# General spin models from noncollinear spin density functional theory and spin-cluster expansion

### v1

@@ II.6 Application to B20-type ${\rm Mn}_{1-x}{\rm Fe}_{x}{\rm Ge}$ and ${\rm Fe}_{1-y}{\rm Co}_{y}{\rm Ge}$
Experimentally, while the helical period $\lambda$ of $\rm FeGe$ exhibits only weak temperature dependence from a quite low temperature to the critical temperature, the spin-stiffness constant $A$ decreases with increasing temperature [40]. Clearly, this implies that an exchange-DMI-only model, with parameters determined at a single temperature, is insufficient to quantitatively describe magnetic order over a broad temperature range. To address this issue, two approaches can be considered: (i) employ a temperature-dependent exchange-DMI-only model or (ii) construct a spin model that includes higher-order interactions capable of describing a broader temperature range (i.e., a broader spin configuration space). However, as we have shown, the $l_{\max}=1$ SCE model cannot adequately renormalize the effects of fluctuations in the local-moment magnitude $|\bm{M}_{\rm Fe}|$; thus, option (ii) is the more appropriate route for capturing magnetism across a wide temperature range. In this context, the advantages of SCE become particularly pronounced: by tuning the model capacity—e.g., increasing $l_{\max}$ and/or the set of included clusters—one can systematically build quantitative spin models. In practice, care must be taken to monitor and mitigate overfitting during model selection. Finally, we note a conceptual distinction from approaches based on the magnetic force theorem (MFT) [49, 50] when evaluating higher-order interactions [51, 52]. MFT extracts interaction parameters from the energy change under infinitesimal spin rotations about a chosen reference state and uses the variation of the single-particle (band) energy within a frozen potential. By contrast, the SCE method samples finite rotations and fits to quantities obtained from fully self-consistent DFT calculations, thereby reflecting the total DFT energy, including band energy and double-counting corrections. Consequently—and provided finite-size effects are adequately controlled—SCE offers a complementary, self-consistent route to estimating higher-order couplings, particularly in regimes with finite rotations or longitudinal fluctuations, and thus serves as a natural counterpart to the MFT analyses.
III Discussion
@@
We have proposed a supercell-based framework for evaluating magnetic interactions of arbitrary form within SCE. In this approach, diverse spin configurations and their site-resolved magnetic torques are computed by constrained noncollinear-spin DFT and used as training data, enabling a highly efficient torque-fitted regression for the SCE model. We applied this framework to representative chiral magnet systems, B20-type $\rm FeGe$ and its variants ${\rm Mn}_{1-x}{\rm Fe}_{x}{\rm Ge}$ and ${\rm Fe}_{1-y}{\rm Co}_{y}{\rm Ge}$. The computed spiralization constant $D$ agrees with other theoretical studies, and the compositions where the helical period diverges are also correctly reproduced. As a distinctive strength of SCE, we have demonstrated the importance of higher-order interactions in moderately disordered magnetic states of $\rm FeGe$. At elevated temperature, the Fe magnetic moments exhibit a wide variance around the modal value. As a result, the lowest-order truncation fails to maintain torque predictive accuracy, whereas including higher-order terms restores agreement. The variance of the local moment strongly depends on the local magnetic disorder (i.e., relative spin angles between nearest-neighbor spins), and higher-order interactions play a compensating role when these relative angles become large. Thus, although the SCE model does not include an explicit degree of freedom for the moment magnitude, spin-configuration-dependent changes in the moment length are implicitly captured through higher-order interactions.
@@
IV Methods
IV.1 Constrained noncollinear-spin DFT
For the constrained noncollinear-spin DFT calculations, we used the VASP package [60, 61] with the projector augmented-wave (PAW) method [62] and the generalized-gradient approximation functional of Perdew, Burke, and Ernzerhof (PBE) [63]. The constrained local moment approach [29] was employed to fix the directions of the atomic magnetic moments, while allowing their magnitudes to relax. Spin–orbit coupling was included self-consistently. The electronic energy-convergence threshold was set to $10^{-7}$ eV/atom. A plane-wave cutoff energy of $300$ eV and a $5\times 5\times 5$ $k$-point grid were used for a $2\times 2\times 2$ cubic supercell (64 atoms). The lattice constant and internal atomic positions were taken from the experimental values of FeGe reported in Ref. [34]. Atomic radii of $1.242$ Å for Fe and $1.161$ Å for Ge were used for the constrained local moment calculations.
IV.2 Conversion from the SCE model to a conventional spin model
@@
$E^{(2)}_{l=1}$ $=4\pi\sum_{i,j,i\neq j}\bm{Z}_{1}(\hat{\bm{e}}_{i})^{\mathsf{T}}\,\mathbf{J}^{(1)}_{ij}\,\bm{Z}_{1}(\hat{\bm{e}}_{j})$ (56)
$=3\sum_{i,j,i\neq j}\hat{\bm{e}}_{i}^{\mathsf{T}}\,\mathbf{J}^{(1)}_{ij}\,\hat{\bm{e}}_{j}\;\equiv\;\sum_{i,j,i\neq j}\hat{\bm{e}}_{i}^{\mathsf{T}}\,\mathcal{J}_{ij}\,\hat{\bm{e}}_{j},$ (57)
@@
where $\mathbb{I}$ is the identity matrix. The antisymmetric exchange $\mathcal{W}_{ij}$ is in one-to-one correspondence with the Dzyaloshinskii–Moriya (DMI) vector $\bm{D}_{ij}$. At the component level,
@@
so that, combining with $\mathcal{W}_{ij}=\tfrac{1}{2}(\mathcal{J}_{ij}-\mathcal{J}_{ij}^{\mathsf{T}})$,

### v2

@@ III.6 Higher-order interactions for thermally disordered states
Experimentally, while the helical period $\lambda$ of $\rm FeGe$ exhibits only weak temperature dependence from a quite low temperature to the critical temperature, the spin-stiffness constant $A$ decreases with increasing temperature [40]. Clearly, this implies that an exchange-DMI-only model, with parameters determined at a single temperature, is insufficient to quantitatively describe magnetic order over a broad temperature range. To address this issue, two approaches can be considered: (i) employ a temperature-dependent exchange-DMI-only model or (ii) construct a spin model that includes higher-order interactions capable of describing a broader temperature range (i.e., a broader spin configuration space). However, as we have shown, the $l_{\max}=1$ SCE model cannot adequately renormalize the effects of fluctuations in the local-moment magnitude $|\mathfrak{m}_{\rm Fe}|$; thus, option (ii) is the more appropriate route for capturing magnetism across a wide temperature range. In this context, the advantages of SCE become particularly pronounced: by tuning the model capacity—e.g., increasing $l_{\max}$ and/or the set of included clusters—one can systematically build quantitative spin models. In practice, care must be taken to monitor and mitigate overfitting during model selection.
IV Conclusion
@@
We have proposed a supercell-based framework for evaluating magnetic interactions of arbitrary form within SCE. In this approach, diverse spin configurations and their site-resolved magnetic torques are computed by constrained noncollinear-spin DFT and used as training data, enabling a highly efficient torque-fitting regression for the SCE model. We applied this framework to representative chiral magnet systems, B20-type $\rm FeGe$ and its variants ${\rm Mn}_{1-x}{\rm Fe}_{x}{\rm Ge}$ and ${\rm Fe}_{1-y}{\rm Co}_{y}{\rm Ge}$. The computed spiralization constant $D$ agrees with other theoretical studies, and the compositions where the helical period diverges are also correctly reproduced. As a distinctive strength of SCE, we have demonstrated the importance of higher-order interactions in moderately disordered magnetic states of $\rm FeGe$. At elevated temperature, the Fe magnetic moments exhibit a wide variance around the modal value. As a result, the lowest-order truncation fails to maintain torque predictive accuracy, whereas including higher-order terms restores agreement. The variance of the local moment strongly depends on the local magnetic disorder (i.e., relative spin angles between nearest-neighbor spins), and higher-order interactions play a compensating role when these relative angles become large. Thus, although the SCE model does not include an explicit degree of freedom for the moment magnitude, spin-configuration-dependent changes in the moment length are implicitly captured through higher-order interactions.
@@
Appendix
Conversion from the SCE model to a conventional spin model
@@
$E^{(2)}_{l=1}$ $=4\pi\sum_{i,j}\bm{Z}_{1}(\hat{\bm{e}}_{i})^{\mathsf{T}}\,\mathbf{J}^{(1)}_{ij}\,\bm{Z}_{1}(\hat{\bm{e}}_{j})$ (55)
$=3\sum_{i,j}\hat{\bm{e}}_{i}^{\mathsf{T}}\,\mathbf{J}^{(1)}_{ij}\,\hat{\bm{e}}_{j}\;\equiv\;\sum_{i,j}\hat{\bm{e}}_{i}^{\mathsf{T}}\,\mathcal{J}_{ij}\,\hat{\bm{e}}_{j},$ (56)
@@
where $\mathbb{I}$ is the identity matrix. The antisymmetric exchange $\mathcal{W}_{ij}$ is in one-to-one correspondence with the DMI vector $\bm{D}_{ij}$. At the component level,
@@
where $\varepsilon_{\alpha\beta\gamma}$ is the Levi–Civita symbol. Combining with $\mathcal{W}_{ij}=\tfrac{1}{2}(\mathcal{J}_{ij}-\mathcal{J}_{ij}^{\mathsf{T}})$, we obtain